OpenStudy (anonymous):

guys can solve this using special product (4x-2y+3z)^2

OpenStudy (anonymous):

so what is the first thing you do with the exponent outside the parenthesis.

OpenStudy (anonymous):

well if you just need the answer carry the exponent by multiplying it by 4,-2,and 3 but because when you multiply whole number's you add the exponents when you divide whole numbers you add the exponents when add whole numbers you multiply the exponents. once doing that you will get your answer.

OpenStudy (anonymous):

I will give you a head start 4*^2= 8 -2*^2 =-4 3*^2=6 so now just add each of the exponents by 2

OpenStudy (anonymous):

and remember x=(x^1)

OpenStudy (anonymous):

^ means to the power of so like and exponent.

OpenStudy (anonymous):

4x-2y +3z * 4x-2y+3z ----------------------------------- 16x^2 -8xy +12zx -8xy +4y^2 -6yz +12zx -6yz +9z^2 __________________________________________ 16x^2-16xy+24zx+4y^2 -12yz +9z^2

OpenStudy (anonymous):

(4x-2y+3z)^2=(4x-2y+3z) *(4x-2y+3z) Actual multiplication is given above.